How do you solve this system by using elimination?
2x-5y=-25
3x+4y=20

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How do you solve this system by using elimination?
2x-5y=-25
3x+4y=20
===============
For elimination, you have to make the coefficients of either x or y match.
---
Multiply the 1st eqn by 3, and the 2nd by 2, then subtract.

Answer by stanbon(75887)

solve this system by using elimination?
2x-5y=-25
3x+4y=20
-----
6x - 15y = 3(-25)
6x + 8y = 2(20)
----------------------
Subtract and solve for "y"::
23y = 40+75
23y = 115
y = 5
------
Solve for "x":
2x - 5(5)= -25
2x = 0
x = 0
----------
Cheers,
Stan H.
----------